OpenStudy (anonymous):

∫sin(60x)((cos(60x))^(1/3))dx

OpenStudy (anonymous):

I JUST WANT THE ANSWER TO COMPARE WITH MINE

OpenStudy (anonymous):

\[=-\frac{1}{80} ( \cos(60x))^{\frac{4}{3}} +C\]